### v2

@@ 4.1 Projection onto the lowest Landau level
To build the setup of the fuzzy sphere, we consider a sphere with radius $R$ and put a $4\pi s$-monopole at its centre. Consider free electrons moving on the sphere. The monopole modifies the single particle Hamiltonian [47, 48, 49, 50]
@@
where $\mu=\theta,\phi$ and the gauge connection is taken as
@@
$Y_{lm}^{(s)}(\hat{\mathbf{n}}),\quad l=s,s+1,\dots,\quad m=-l,\dots,l-1,l,$ (8)
@@
where $\hat{\mathbf{n}}$ is the unit vector of the point on the sphere specified by angular coordinates $\theta$ and $\phi$, and the energies are
$E_{l}=\frac{1}{2MR^{2}}(l(l+1)-s^{2}).$ (9)
@@
Each level, known as a Landau level, has a degeneracy $(2l+1)$. Specifically, the wavefunctions on the lowest Landau level (LLL) $l=s$ are easy to write out :
$Y_{sm}^{(s)}(\hat{\mathbf{n}})=C_{m}e^{im\phi}\cos^{s+m}\frac{\theta}{2}\sin^{s-m}\frac{\theta}{2},\quad C_{m}=\frac{1}{{\sqrt{4\pi\mathrm{B}(s+m+1,s-m+1)}}},$ (10)
@@
where $C_{m}$ is the normalising factor, and $\mathrm{B}$ is the Euler’s beta function. The LLL has a degeneracy $N_{m}=2s+1$.
@@
We now consider $N_{f}$ flavours of fermions moving on the sphere, characterised by the second-quantised fermion operator $\hat{\psi}_{f}(\hat{\mathbf{n}})$, with a flavour index $f=1,\dots,N_{f}$. We partially fill the lowest Landau level and set the single energy gap to be much larger than the scale of interaction $H_{0}\gg H_{\textrm{int}}$, so that the quantum fluctuation can be constrained on the lowest Landau level. In practice, we often fill integer number of flavours $N_{e}=kN_{m}$ ($k\in\mathbb{Z}$) so that a quantum Hall ferromagnet (i.e., the state where integer number of LLLs are fully filled) is preferred in the absence of interaction, for which the charge degree of freedom is gapped and does not couple strongly to the gapless CFT degree of freedom when the interactions are introduced.
We then project the system onto the LLL. Technically, this can be done by writing the fermion operators in terms of the annihilation operators of the LLL orbitals
$\hat{\psi}_{f}(\hat{\mathbf{n}})=\sum_{m=-s}^{s}Y^{(s)}_{sm}(\hat{\mathbf{n}})\hat{c}_{mf},$ (11)
@@
where $\hat{c}^{(\dagger)}_{mf}$ annihilates/creates an electron with $L^{z}$-quantum number $m$ at the $f$-th flavour of the lowest Landau level. Hereafter, we will omit the hats on the operators.
@@
After the projection, we obtain a finite Hilbert space on which numerical simulations can be carried out. For this purpose, the system is analogous to a length-$(2s+1)$ spin chain with long-range interaction, where different Landau level orbitals behave like the lattice sites. The difference is that the $(2s+1)$ orbital forms a spin-$s$ representation of the $\mathrm{SO}(3)$ rotation group, and in this way the continuous rotation symmetry is preserved. The exact rotation symmetry shortens the RG flow from the UV to the IR and reduces the finite-size effect, so that the numerical results are considerably accurate even at a small system size.
@@
The word ‘fuzzy’ means non-commutativity [46]. Here, the magnetic field results in the non-commutativity of the coordinates. More concretely, we write the coordinate operators as a matrix on the lowest Landau level
$X^{\mu}_{m_{1}m_{2}}=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,n^{\mu}\bar{Y}_{sm_{1}}^{(s)}(\hat{\mathbf{n}})Y_{sm_{2}}^{(s)}(\hat{\mathbf{n}}).$ (12)
@@
$\mathbf{X}_{\mu}\mathbf{X}^{\mu}=\frac{s}{s+1}\mathbb{I},\qquad[\mathbf{X}^{\mu},\mathbf{X}^{\nu}]=\frac{1}{s+1}i\epsilon^{\mu\nu\rho}\mathbf{X}_{\rho}.$ (13)
The first equation involves the radius $R$ of the sphere, and the second equation involves the magnetic length $l_{B}$ that determines the non-commutativity.
$\mathbf{X}_{\mu}\mathbf{X}^{\mu}=R^{2}\mathbb{I},\qquad[\mathbf{X}^{\mu},\mathbf{X}^{\nu}]=l_{B}^{2}\,i\epsilon^{\mu\nu\rho}(\mathbf{X}_{\rho}/R).$
We can take $l_{B}=1$ as the unit length. In this way, the radius scales with the square root of the number of orbitals
$R/l_{B}=[s(s+1)]^{1/4}\sim\sqrt{N_{m}}.$ (14)
The thermodynamic limit can be taken as $N_{m}\to\infty$, where a regular sphere is recovered.
@@ 4.2 Density operator
$n_{M}(\hat{\mathbf{n}})=\psi_{f^{\prime}}^{\dagger}(\hat{\mathbf{n}})M_{f^{\prime}f}\psi_{f}(\hat{\mathbf{n}}).$ (15)
Here, the matrix insertion $M$ puts the density operators in a certain representation of the flavour symmetry. For example, for a 2-flavour system, $M$ can be taken as the Pauli matrices $\mathbb{I},\sigma^{x},\sigma^{y},\sigma^{z}$ ; for a system with $N_{f}$ flavours in the fundamental representation of $\mathrm{SU}(N_{f})$ flavour symmetry, one can put $n_{M}$ in the singlet or adjoint representation
$n_{S}(\hat{\mathbf{n}})$ $=\psi_{c}^{\dagger}(\hat{\mathbf{n}})\psi^{c}(\hat{\mathbf{n}})$
$n_{a}{}^{b}(\hat{\mathbf{n}})$ $=\psi_{a}^{\dagger}(\hat{\mathbf{n}})\psi^{b}(\hat{\mathbf{n}})-\tfrac{1}{N}\delta_{a}{}^{b}\psi_{c}^{\dagger}(\hat{\mathbf{n}})\psi^{c}(\hat{\mathbf{n}}).$ (16)
Like the fermion operator, the density operator can also be expressed in the orbital space [^4] Under this convention, the action of $n_{M,lm}$ will decrease the angular momentum $L^{z}$ by $m$. Our convention of $n_{lm}$ differs from Ref. [2] by a factor of $R^{2}$. We take this convension in alignment with the implementation in the code.
$n_{M}(\hat{\mathbf{n}})=\sum_{lm}Y_{lm}(\hat{\mathbf{n}})n_{M,lm}.$ (17)
@@
$n_{M,lm}$ $=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{lm}(\hat{\mathbf{n}})n_{M}(\hat{\mathbf{n}})$
$=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{lm}(\hat{\mathbf{n}})\left(\sum_{m_{1}}\bar{Y}^{(s)}_{sm_{1}}(\hat{\mathbf{n}})c^{\dagger}_{m_{1}f_{1}}\right)M_{f_{1}f_{2}}\left(\sum_{m_{2}}Y^{(s)}_{sm_{2}}(\hat{\mathbf{n}})c_{m_{1}f_{2}}\right)$
$=\sum_{m_{1}m_{2}}c^{\dagger}_{m_{1}f_{1}}M_{f_{1}f_{2}}c_{m_{1}f_{2}}\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{lm}(\hat{\mathbf{n}})\bar{Y}^{(s)}_{sm_{1}}(\hat{\mathbf{n}})Y^{(s)}_{sm_{2}}(\hat{\mathbf{n}})$
$=\sum_{m_{1}}c^{\dagger}_{m_{1}f_{1}}M_{f_{1}f_{2}}c_{m+m_{1},f_{2}}\times$
$\qquad\qquad(-1)^{s+m+m_{1}}(2s+1)\sqrt{\frac{2l+1}{4\pi}}\begin{pmatrix}s&l&s\\ m_{1}&m&-m_{1}-m\end{pmatrix}\begin{pmatrix}s&l&s\\ s&0&-s\end{pmatrix}.$ (18)
@@
where ($\ \bullet&\bullet&\bullet\end{smallmatrix}\right)$ is the $3j$-symbol [96]. In this way, we have fully expressed the density operator in terms of the operators in the orbital space $c^{(\dagger)}_{mf}$.
@@ 4.3 Density-density interaction
The most straightforward way to construct an interaction term is to add a density-density interaction with a potential function [^5] This is, however, not the simplest construction and we will present the simpler construction in terms of pseudopotentials in the next section.
$H_{\mathrm{int}}=\int\mathrm{d}^{2}\hat{\mathbf{n}}_{1}\,\mathrm{d}^{2}\hat{\mathbf{n}}_{2}\,U(|\hat{\mathbf{n}}_{1}-\hat{\mathbf{n}}_{2}|)n_{M}(\hat{\mathbf{n}}_{1})n_{M}(\hat{\mathbf{n}}_{2}).$ (20)
@@
where $\mathbf{r}_{12}=\hat{\mathbf{n}}_{1}-\hat{\mathbf{n}}_{2}$ and $|\mathbf{r}_{12}|=2\sin\theta_{12}/2$. Conversely
@@
Specifically, for local and super-local interactions [^6] Usually, to keep the interaction Hamiltonian extensive, the factors scale with system size as $g_{0}\propto R^{-2}$ and $g_{1}\propto R^{-4}$
$U(|\mathbf{r}_{12}|)$ $=g_{0}\delta(\mathbf{r}_{12}),$ $\tilde{U}_{l}$ $=g_{0}(2l+1)$
$U(|\mathbf{r}_{12}|)$ $=g_{1}\nabla^{2}\delta(\mathbf{r}_{12}),$ $\tilde{U}_{l}$ $=-g_{1}l(l+1)(2l+1).$ (23)
By expanding the density operators into the orbital space and completing the integrals,
$H_{\textrm{int}}=\sum_{lm}\frac{4\pi\tilde{U}_{l}}{2l+1}n^{\dagger}_{M,lm}n_{M,lm}.$ (24)
@@
$-h\int\mathrm{d}^{2}\hat{\mathbf{n}}\,n_{x}(\hat{\mathbf{n}})$
@@
$\int\mathrm{d}^{2}\hat{\mathbf{n}}_{1}\,\mathrm{d}^{2}\hat{\mathbf{n}}_{2}\,U(|\mathbf{r}_{12}|)n_{\uparrow}(\hat{\mathbf{n}}_{1})n_{\downarrow}(\hat{\mathbf{n}}_{2}),$
@@
$n_{x}(\hat{\mathbf{n}})=\psi^{\dagger}_{\downarrow}(\hat{\mathbf{n}})\psi_{\uparrow}(\hat{\mathbf{n}})+\psi^{\dagger}_{\uparrow}(\hat{\mathbf{n}})\psi_{\downarrow}(\hat{\mathbf{n}}),\quad n_{\uparrow}(\hat{\mathbf{n}})=\psi^{\dagger}_{\uparrow}(\hat{\mathbf{n}})\psi_{\uparrow}(\hat{\mathbf{n}}),\quad n_{\downarrow}(\hat{\mathbf{n}})=\psi^{\dagger}_{\downarrow}(\hat{\mathbf{n}})\psi_{\downarrow}(\hat{\mathbf{n}}),$
@@
$H_{\textrm{int}}=\int\mathrm{d}^{2}\hat{\mathbf{n}}_{1}\,\mathrm{d}^{2}\hat{\mathbf{n}}_{2}\,U(|\mathbf{r}_{12}|)n_{\uparrow}(\hat{\mathbf{n}}_{1})n_{\downarrow}(\hat{\mathbf{n}}_{2})-h\int\mathrm{d}^{2}\hat{\mathbf{n}}\,n_{x}(\hat{\mathbf{n}}).$ (25)
@@
4.4 Interaction in terms of pseudopotentials
Another way that is much more convenient to construct the four-fermion interaction terms is through Haldane pseudopotential [47, 97]. To explain the idea, we take the 3d Ising model as an example. We first classify all the fermion bilinears $\lambda_{m_{1}f_{1}m_{2}f_{2}}c_{m_{1}f_{1}}c_{m_{2}f_{2}}$. To simplify the discussion, we can take a specific flavour index $\lambda_{m_{1}m_{2}}c_{m_{1}\uparrow}c_{m_{2}\downarrow}$. The fermion bilinears can be classified into irreducible representations (irreps) of $\mathrm{SO}(3)$ rotation symmetry. Since $c_{mf}$ carries the spin-$s$ representation, the spin of its bilinear ranges from 00 to $2s$ and takes integer values. The spin-$(2s-l)$ combination reads

### v3

@@ 4.1 Projection onto the Lowest Landau Level
To build the set-up of the fuzzy sphere, we consider a sphere with radius $R$ and put a $4\pi s$-monopole at its centre. Consider free electrons moving on the sphere. The monopole modifies the single particle Hamiltonian [63, 64, 65, 66]
@@
where $\mu=\theta,\phi$ and the gauge connexion is taken as
@@
$\Phi_{nm}(\mathbf{r})=\frac{1}{R}Y_{lm}^{(s)}(\hat{\mathbf{n}}),\qquad n=0,1,\dots,\quad l=n+s,\quad m=-l,\dots,l-1,l,$ (7)
@@
where $\hat{\mathbf{n}}$ is the unit vector of the point on the sphere specified by angular co-ordinates $\theta$ and $\phi$, and the energies are
$E_{n}=\frac{1}{2MR^{2}}n(2s+n+1).$ (8)
@@
Each level, known as a Landau level, has a degeneracy $(2l+1)$. Specifically, the wave-functions on the lowest Landau level (LLL) $n=0,l=s$ are easy to write out :
$\Phi_{0m}(\mathbf{r})=\frac{1}{R}Y_{sm}^{(s)}(\hat{\mathbf{n}}),\qquad Y_{sm}^{(s)}(\hat{\mathbf{n}})=C_{m}e^{im\phi}\cos^{s+m}\frac{\theta}{2}\sin^{s-m}\frac{\theta}{2}$ (9)
@@
where $C_{m}=1/\sqrt{4\pi\mathrm{B}(s+m+1,s-m+1)}$ is the normalising factor, and $\mathrm{B}$ is the Euler’s beta function. The LLL has a degeneracy $N_{m}=2s+1$.
@@
We now consider $N_{f}$ flavours of fermions moving on the sphere, characterised by the second-quantised fermion operator $\psi_{f}(\mathbf{r})$, with a flavour index $f=1,\dots,N_{f}$. We partially fill the lowest Landau level and set the single energy gap to be much larger than the scale of interaction $H_{0}\gg H_{\text{int}}$, so that the quantum fluctuation can be constrained on the lowest Landau level. In practice, we often fill integer number of flavours $N_{e}=kN_{m}$ ($k\in\mathbb{Z}$) so that a quantum Hall ferromagnet (i. e. the state where integer number of LLLs are fully filled) is preferred in the absence of interaction, for which the charge degree of freedom is gapped and decouples from the gapless CFT degree of freedom when the interactions are introduced. [^5] The charge degree of freedom cannot flow to any gapless CFT local operator because the charged operators carry projective representations of the rotation group due to the non-commutativity while the CFT operators carry linear prepresentations.
We then project the system onto the LLL. Technically, this can be done by writing the fermion operators in terms of the annihilation operators of the LLL orbitals [^6] In the old version, we used a different convension $\psi_{f}(\hat{\mathbf{n}})_{\text{(old)}}=\sum_{m=-s}^{s}Y^{(s)}_{sm}(\hat{\mathbf{n}})c_{mf}.$ In the old convension, the components of density operator $n_{M,lm}$ are extensive and their action decreases the angular momentum $L^{z}$ by $m$ ; in the current convension, $n_{M,lm}$ is intensive and its action increases the angular momentum $L^{z}$ by $m$. In the code, two parameters ObsNormRadSq and ObsMomIncr control the convension.
$\psi^{\dagger}_{f}(\mathbf{r})=\sum_{m=-s}^{s}\Phi_{0m}(\mathbf{r})c^{\dagger}_{mf}=\frac{1}{R}\sum_{m=-s}^{s}Y^{(s)}_{sm}(\hat{\mathbf{n}})c^{\dagger}_{mf},$ (10)
@@
where $c^{(\dagger)}_{mf}$ annihilates/creates an electron with $L^{z}$-quantum number $m$ at the $f$-th flavour of the lowest Landau level. Hereafter, we will omit the hats on the operators.
@@
After the projection, we obtain a finite Hilbert space on which numerical simulations can be carried out. For this purpose, the system is analogous to a length-$(2s+1)$ spin chain with long-range interaction, where different Landau level orbitals behave like the lattice sites. The difference is that the $(2s+1)$ orbitals form a spin-$s$ representation of the $\mathrm{SO}(3)$ rotation group, and in this way the continuous rotation symmetry is preserved. The exact rotation symmetry shortens the RG flow from the UV to the IR and reduces the finite-size effect, so that the numerical results are considerably accurate even at a small system size.
@@
The word “fuzzy” means non-commutativity [62]. Here, the magnetic field results in the non-commutativity of the co-ordinates. More concretely, we write the co-ordinate operators as a matrix on the lowest Landau level
$X^{\mu}_{m_{1}m_{2}}=\int\mathrm{d}^{2}\mathbf{r}\,x^{\mu}\frac{Y_{sm_{1}}^{(s)}(\mathbf{r})}{R}\frac{\bar{Y}_{sm_{2}}^{(s)}(\mathbf{r})}{R}.$ (11)
@@
$\mathbf{X}_{\mu}\mathbf{X}^{\mu}=\frac{s}{s+1}R^{2}\mathbb{I},\qquad[\mathbf{X}^{\mu},\mathbf{X}^{\nu}]=\frac{1}{s+1}i\epsilon^{\mu\nu\rho}R\mathbf{X}_{\rho}.$ (12)
The first equation involves a renormalised radius $\tilde{R}$ of the sphere, and the second equation involves the magnetic length $l_{B}$ that determines the non-commutativity.
$\mathbf{X}_{\mu}\mathbf{X}^{\mu}=\tilde{R}^{2}\mathbb{I},\qquad[\mathbf{X}^{\mu},\mathbf{X}^{\nu}]=l_{B}^{2}\,i\epsilon^{\mu\nu\rho}(\mathbf{X}_{\rho}/R).$
Their ratio scales with the square root of the number of orbitals
$\tilde{R}/l_{B}=[s(s+1)]^{1/4}\sim\sqrt{N_{m}}.$ (13)
We can take $l_{B}=1$ as the unit length. In this way, the thermodynamic limit can be taken as $N_{m}\to\infty$, where a regular sphere is recovered. Hereafter we take the radius of the sphere $R=\sqrt{N_{m}}$
@@ 4.2 Density Operator
$n_{M}(\mathbf{r})=\psi_{f^{\prime}}^{\dagger}(\mathbf{r})M_{f^{\prime}f}\psi_{f}(\mathbf{r}).$ (14)
Here, the matrix insertion $M$ puts the density operators in a certain representation of the flavour symmetry. E. g., for a 2-flavour system, $M$ can be taken as the Pauli matrices $\mathbb{I},\sigma^{x},\sigma^{y},\sigma^{z}$ ; for a system with $N_{f}$ flavours in the fundamental representation of $\mathrm{SU}(N_{f})$ flavour symmetry, one can put $n_{M}$ in the singlet or adjoint representation
$n_{S}(\mathbf{r})$ $=\psi_{c}^{\dagger}(\mathbf{r})\psi^{c}(\mathbf{r})$
[…]
Like the fermion operator, the density operator can also be expressed in the orbital space
$n_{M}(\mathbf{r})=\sum_{lm}Y_{lm}(\hat{\mathbf{n}})n_{M,lm}.$ (16)
@@
$n_{M,lm}$ $=\frac{1}{R^{2}}\int\mathrm{d}^{2}\mathbf{r}\,\bar{Y}_{lm}(\hat{\mathbf{n}})n_{M}(\mathbf{r})$
$=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{lm}(\hat{\mathbf{n}})\left(\frac{1}{R}\sum_{m_{1}}Y^{(s)}_{sm_{1}}(\hat{\mathbf{n}})c^{\dagger}_{m_{1}f_{1}}\right)M_{f_{1}f_{2}}\left(\frac{1}{R}\sum_{m_{2}}\bar{Y}^{(s)}_{sm_{2}}(\hat{\mathbf{n}})c_{m_{1}f_{2}}\right)$
[…]
$=\frac{1}{R^{2}}\sum_{m_{1}}c^{\dagger}_{m_{1}f_{1}}M_{f_{1}f_{2}}c_{m-m_{1},f_{2}}$
$\qquad\qquad\times(-1)^{s+2m-m_{1}}(2s+1)\sqrt{\frac{2l+1}{4\pi}}\begin{pmatrix}l&s&s\\ -m&m_{1}&-m_{1}+m\end{pmatrix}\begin{pmatrix}l&s&s\\ 0&-s&s\end{pmatrix}$. (17)
@@
[…]
@@ 4.3 Density-Density Interaction
The most straightforward way to construct an interaction term is to add a density-density interaction with a potential function [^7] Computationally, this is not the simplest construction and we will present the simpler construction in terms of pseudo-potentials in the next section.
$H_{\text{int}}=\int\mathrm{d}^{2}\mathbf{r}_{1}\,\mathrm{d}^{2}\mathbf{r}_{2}\,U(|\mathbf{r}_{1}-\mathbf{r}_{2}|)n_{M}(\mathbf{r}_{1})n_{M}(\mathbf{r}_{2}).$ (19)
@@
where the cordial distance $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$ and $|\mathbf{r}_{12}|=2R\sin\theta_{12}/2$. Conversely
@@
Specifically, for local and super-local interactions
$U(|\mathbf{r}_{12}|)$ $=g_{0}\delta(\mathbf{r}_{12}),$ $\tilde{U}_{l}$ $=\frac{g_{0}}{R^{2}}(2l+1)$
$U(|\mathbf{r}_{12}|)$ $=g_{1}\nabla^{2}\delta(\mathbf{r}_{12}),$ $\tilde{U}_{l}$ $=-\frac{g_{1}}{R^{4}}l(l+1)(2l+1).$ (22)
Here we make use of the conversion relations
$\delta(\mathbf{r})=\frac{1}{R^{2}}\delta(\hat{\mathbf{n}}),\qquad\nabla^{2}_{\mathbf{r}}=\frac{1}{R^{2}}\nabla^{2}_{\hat{\mathbf{n}}}$
for $\mathbf{r}=R\hat{\mathbf{n}}$. By expanding the density operators into the orbital space and completing the integrals,
$H_{\text{int}}=\sum_{lm}\frac{4\pi R^{4}\tilde{U}_{l}}{2l+1}n^{\dagger}_{M,lm}n_{M,lm}.$ (23)
@@
$h\int\mathrm{d}^{2}\mathbf{r}\,n_{x}(\mathbf{r})$
@@
$\int\mathrm{d}^{2}\mathbf{r}_{1}\,\mathrm{d}^{2}\mathbf{r}_{2}\,U(|\mathbf{r}_{12}|)n_{\uparrow}(\mathbf{r}_{1})n_{\downarrow}(\mathbf{r}_{2}),$
@@
$n_{x}(\mathbf{r})=\psi^{\dagger}_{\downarrow}(\mathbf{r})\psi_{\uparrow}(\mathbf{r})+\psi^{\dagger}_{\uparrow}(\mathbf{r})\psi_{\downarrow}(\mathbf{r}),\quad n_{\uparrow}(\mathbf{r})=\psi^{\dagger}_{\uparrow}(\mathbf{r})\psi_{\uparrow}(\mathbf{r}),\quad n_{\downarrow}(\mathbf{r})=\psi^{\dagger}_{\downarrow}(\mathbf{r})\psi_{\downarrow}(\mathbf{r}),$
@@
$H_{\text{int}}=\int\mathrm{d}^{2}\mathbf{r}_{1}\,\mathrm{d}^{2}\mathbf{r}_{2}\,U(|\mathbf{r}_{12}|)n_{\uparrow}(\mathbf{r}_{1})n_{\downarrow}(\mathbf{r}_{2})-h\int\mathrm{d}^{2}\mathbf{r}\,n_{x}(\mathbf{r}).$ (24)
@@
4.4 Interaction in Terms of Pseudo-Potentials
Another way that is computationally more convenient to construct the four-fermion interaction terms is through Haldane pseudo-potential [63, 141]. To explain the idea, we take the 3D Ising model as an example. We first classify all the fermion bilinears $\lambda_{m_{1}f_{1}m_{2}f_{2}}c_{m_{1}f_{1}}c_{m_{2}f_{2}}$. To simplify the discussion, we can take a specific flavour index $\lambda_{m_{1}m_{2}}c_{m_{1}\uparrow}c_{m_{2}\downarrow}$. The fermion bilinears can be classified into irreducible representations (irreps) of $\mathrm{SO}(3)$ rotation symmetry. Since $c_{mf}$ carries the spin-$s$ representation, the spin of its bilinear ranges from 0 to $2s$ and takes integer values. The spin-(2sl)(2s-l) combination reads